The Distributive Property (Fractions)
1/3 (6x-9)+x= 15

Answer by BGluvsmath(64)

1/3 (6x-9)+x= 15
First, distribute: 2x-3+x=15
Simplify: 3x-3=15
Add 3 to both sides: 3x=18
divide both sides by 3 and x=6